OpenStudy (anonymous):

A rock weighs 33.6 N on Planet X and 49 N on Earth. What is g on Planet X? 6.72 m/sec2 168 m/sec2 0.68 m/sec2 672 m/sec2

OpenStudy (anonymous):

@cutiepie15

OpenStudy (anonymous):

I am in 7th what grade is this?

OpenStudy (anonymous):

i am in 12th grade but i need help with physical science its for 9th graders i just have never taken this class i have taken other classes but this is not familiar to me do you you know someone that can help me

OpenStudy (anonymous):

@Tootles143 @TwilightRain1 i need help i just never needed to take this

OpenStudy (anonymous):

@SaltLife23 @ckellar2

OpenStudy (anonymous):

maybe b?

OpenStudy (anonymous):

Weight is directly proportional to g Because \[W = mg\] Hence, \[\frac{g_X}{g_{earth}} = \frac{W_x}{W_{earth}}\]

OpenStudy (anonymous):

We are weighing the same rock, just different locations. On Planet X, the rock weighs 69% of what it does on earth. (33.6/49=.68571429, which roughly rounds to 69%). If the acceleration of gravity on earth is 9.8 m/s^2, then we just need to find what roughly 69% of that is. (9.8x.68571429=6.72; Planet X has a gravity of 6.72 earth gravities)

OpenStudy (anonymous):

so its the first one? 6.72?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

okay thank you